A student averages 84 on five tests. If the student received grades of 84,82,81, and 79 on the first four exams, what grade did the student receive on the fifth exam?
Answer by checkley77(12844)

84,82,81, and 79
(84+82+81+79+x)/5=84
(326+x)/5=84 cross multiply
326+x=5*84
326+x=420
x=420-326
x=94 ans. for the grade of the fifth test.
Proof:
(84+82+81+79+94)/5=84
420/5=84
84=84